OpenStudy (anonymous):

What is 8/3 divide by 3?

OpenStudy (freckles):

\[\frac{8}{3} \div \frac{3}{1} =\frac{8}{3} \cdot \frac{1}{3}\] multiply fractions by multiplying straight across on top and then on bottom